OpenStudy (anonymous):

Please help...I forget how to do this: Find all Values of x, in radians, such that 2cos^2x+5cosx -3 =0 where 0 is less than or equel to x which is less than or equel to pi.

OpenStudy (anonymous):

Can I use Quadratic equation some how?

OpenStudy (raden):

factor out, it can be 2cos^2x+5cosx -3 =0 (2cosx -1)(cosx + 3) = 0 for the zeroes satisfied : 2cosx -1 = 0 or cosx + 3 = 0 continue it ...

OpenStudy (anonymous):

2cosx=1 cosx=1/2?

OpenStudy (anonymous):

cosx+3 = 0 cosx = -3

OpenStudy (raden):

u knowed that the minimum value for cos x = -1, so there is no solution for cosx=-3. just find the solution of cosx = 1/2

OpenStudy (raden):

what angle the cos has value 1/2 ?

OpenStudy (anonymous):

60degrees

OpenStudy (anonymous):

and 300, but can't be 300 because its 0-pi

OpenStudy (anonymous):

Right? Thanks mate! Your a good teacher

OpenStudy (raden):

yes.. u r right. dont forget convert 60 deg to rad :)

OpenStudy (anonymous):

1.047 ;) thanks again

OpenStudy (raden):

60 degrees = 60/180 pi = pi/3